### v1

# MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels

Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), which effectively assesses both uni-modal and omni-modal understanding capabilities. The benchmark consists of 1880 human curated samples, across 44 task types, and a innovative multi-step open-ended question type that better assess complex reasoning tasks. Experimental result shows the compositional law between cross-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels
Chen Chenthanks: Equal contribution.   ZeYang Hu11footnotemark: 1   Fengjiao Chenthanks: Project Lead. Liya Ma   Jiaxing Liu   Xiaoyu Li22footnotemark: 2   Xuezhi Cao Meituan, Beijing, China. {chenchen165, huzeyang, chenfengjiao02, maliya06}@meituan.com {liujiaxing10, lixiaoyu28, caoxuezhi}@meituan.com
1 Introduction
Multimodal artificial intelligence has undergone extensive researches in visual language model and audio language model, with current advancements progressing toward unifying visual, audio and language modalities, collectively termed omni models. The evaluation paradigm for these models has consequently expanded from assessing uni-modal understanding capabilities (i.e. visual understanding, audio understanding) to the next-level of intelligence (i.e. omni-modal understanding).
Existing omni-model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
The ideal omni-model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without video or audio, and 25% questions from OmniBench contain erroneous answers. These designs limit the evaluation and analysis of omni models’ abilities.
Addressing theses limitations, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
Main Contributions:
1. Propose the first MultiModal All in One Benchmark evaluation framework specifically designed for omni models, which efficiently assesses both uni-modal and omni-modal understanding capabilities. This framework verifies the compositional law between omni-modal and uni-modal capability. For weaker models, this manifests itself as a "short-board effect" (where performance is constrained by the weakest modality), while stronger models demonstrate emergent ability across modalities.
2. Establish a systematic and unified dataset framework, featuring complete manual annotation to ensure high quality and diversity. This approach effectively prevents data contamination while better aligning with practical application scenarios. As a result, the MMAO-Bench comprises 44 different types of tasks and is 100% passing the quality check, which significantly surpassing existing datasets.
3. Beyond conventional multiple-choice formats, the evaluation incorporates innovative Multi-Step Open-Ended Question tasks to show a more realistic and discriminative evaluation result on complex reasoning.
2 Related Work
2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024; Wang et al., 2024a, b; Liu et al., 2024c; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024b; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024c) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while qwen-3-omni(Xu et al., 2025b) provides leading performance in open-sourced models. However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
Due to the limited quality and coverage of existing benchmarks, we introduce a new benchmark MMAO-Bench which comprise 1250 human crafted questions for omni-modal and 630 enhanced samples for uni-modal, across 44 task types with 98% cross-modality solvable. The materials are collected from in-house dataset with manual annotation which prevents data contamination while better aligning with practical application scenarios. Beyond conventional multiple-choice formats, the evaluation incorporates innovative multi-step open-ended question tasks to show a more realistic and discriminative evaluation result on complex reasoning. Besides the self-constructed omni-modal dataset, we construct a uni-modal dataset by aggregating public uni-model datasets systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency. In this way, our benchmark can have a thoroughly evaluation that requires omni models to keep uni-modal abilities and obtain omni-modal abilities at the same time.
Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
MMBench I - - 80%private 20 3217 MC EN/ZH
MMAU A - - 15%private 27 10000 MC EN
MVBench V - - public 20 4000 MC EN
OmniBench I+A 75% 90% public 8 1142 MC EN
AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
WorldSense V+A 99% 23% public 26 3172 MC EN
Daily-Omni V+A - - public 6 1197 MC EN
MMAO-Bench I+V+A I/V/A 100% 98% 90%private 44 1880 MC/MO EN/ZH
Table 1: Comparison of MultiModal Benchmarks. Where I,A,V and T indicate image, audio, video and text modality. The accuracy of question answer pairs and the percent of questions needing omni-modal to solve is shown in Acc. and Solvable. Source indicates where the materials are collected from, public datasets or private datasets that can prevent data contamination. For QA type, MC, SO and MO indicates multi-choise QA, single-open question QA and multi-open question QA. EN and ZH indicates english and chinese.
3 Method
In this section, we first introduce the comprehensive design of model ability system in Section.3.1. The the high-quality data construction pipeline are introduced in Section.3.2. The multi-step open-ended questions are introduced in Section.3.4. Finally, to improve the evaluation efficiency, we introduce a general evaluation dataset compression method in Section.3.3.2.
3.1 Model Ability System
As shown in Figure.1, the model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
Perception Layer Structured through six recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross- Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
Reasoning Layer extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning, Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
This ability system combines uni-modal and cross-modal abilities which enables cross-modal measurement that is particularly critical for advanced omni-models. Based on this system, we create a diversity dataset with 44 task types distributed as shown in Figure.2.
Figure 1: Illustration of the omni-modal ability system proposed in MMAO-Bench.
Figure 2: Data distribution in model abilities and corresponding task types.
3.2 Omni-modal Dataset Construction
We have established a standardized and efficient data construction process (Figure.3) that produces high-quality and high-diversity dataset.
Figure 3: Data Construction Pipeline
Preparation: (1) Task Design. Deconstruct cross-modal capability systems and task types. Details can be seen in Figure.2. (2) Material Collection. Our materials have the following three characteristics: a. Diverse Sources. The majority of our materials are real-world photos and videos collected through crowdsourcing, with another portion sourced from copyright-free website. Additionally, a small fraction comes from high-quality public datasets such as MMVU, LongVideoBench, and VideoVista. b. Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on. c. Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world. (3) Material filtering. Eliminate meaningless, illogical, and low-quality materials, and categorize the remaining materials by theme to create a material library. Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
QA Annotation: (1) Annotators. It mainly includes two parts: human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity. (2) Annotation process. First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags. Second, following data construction requirements, they then design prompts and corresponding answers. Third, to enhance data authenticity, all dialogue audio is recorded manually. Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
Quality Inspection: We have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Secondly, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unanswerable or ambiguous after removing any one modality, it proves the three-modality solvability of the data. Finally, manual quality inspection and revision are performed.
3.3 Uni-modal Dataset Improvement
3.3.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue. To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench. As shown in Figure4, experiments show that the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
Figure 4: Comparison result of privatization improvement. After improvement, the performances among models are more progressively and distinguishable.
In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey and WorldSense, accounting for 11% of the total). The specific selection logic is as follows:
1. Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting video + audio modality combinations with a lower self-construction proportion for supplementation, followed by image + audio.
2. Diversity: Supplement material types not covered in self-construction data to enhance diversity.
3. High Quality: Pay attention to the quality of datasets during screening (whether uni-modal answers are possible, accuracy, reasonableness, etc.). For most world senses that can be answered with uni-modal data, they will not be included.
4. Discriminative: Pay attention to the performance of this data subset on the model during screening, and remove overly difficult subsets with little discriminative power.
3.3.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling method. This method achieves efficient compression and high evaluation consistency through the following steps:
Question Characterization: Represent each question as an x-dimensional vector, where dimensions correspond to scores from different models on that question.
Cluster-based Stratification: Utilize the K-Means algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
Hierarchical Sampling: Determine the sample size for each stratum based on cluster size proportions, and construct the final evaluation subset through simple random sampling.
Validity Verification: Use Margin of Error (MoE) and Root Mean Square Error (RMSE) to measure theoretical and practical errors respectively. Verify the consistency between model rankings and scores using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
Finally, in order to ensure statistical stability, we repeat above steps by employing 5 random splits and 10-fold cross-validation. This method determines the optimal sample size through cost-benefit curve analysis, reducing evaluation costs by over 90% while maintaining accuracy (Figure.8).
3.4 Multi-Step Open-Ended Questions
Cross-modality problem requires the model to understand among multiple modalities with multi-step reasoning. However, existing evaluation benchmarks primarily focus on isolated problems where questions and answers remain independent of each other. Such a single-step evaluation paradigm cannot effectively assess the ability of a model to understand the complex, cross modality tasks that require multiple logical steps. Besides, most benchmarks adopt multiple-choice questions to verify the model’s performance, which reduces the difficulty of questions (e.g. allows for guessing) and may underestimate the model’s deficiencies in complex tasks. To address these issues, we introduce a new evaluation method, Multi-Step Open-Ended Questions (MO), which breaks down complex questions into multiple interdependent sub-tasks and requires the model to solve them step by step in the form of open-ended text. The construction process and example are shown in Figure.5. The multi-step open-ended question is worth 10 points. During the process of providing a reference answer, scores are assigned to each sub-question based on its difficulty and necessity for answering the final question, which facilitates the evaluation of the model’s accuracy at each step.
Experiments (4.4.1) show that compared with single-step evaluation method (e.g. multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
Figure 5: Example of multi-step open-ended questions.
Table 2: Cross-modal benchmark results on multimodal models
Model Visual Audio Omni-MC Omni-MO
MiniCPM-o-2.6 41.85 63.10 26.77 21.92
Qwen-2.5-omni-3B 45.71 65.81 27.26 25.72
Baichuan-omni-1.5 49.25 66.50 30.32 20.48
Ming-lite-omni-1.5 48.68 71.70 32.10 26.08
Qwen-2.5-omni-7B 52.81 68.11 33.23 29.12
Gemini-2.0-Flash 66.80 74.41 46.26 40.96
Gemini-2.5-Flash 76.64 75.20 68.23 62.00
Gemini-2.5-Pro 80.10 78.70 76.61 68.84
Table 3: Cross-modal analysis on model ability system
Model Perception Reasoning Overall
Cross-modal
Alignment
Cross-modal
Recognition
Overall
Sequential
Reasoning
Spatial
Reasoning
General
Reasoning
Overall
MiniCPM-o-2.6 25.55 36.11 27.75 38.78 22.94 26.36 26.40 26.77
Qwen-2.5-omni-3B 27.74 50.00 32.37 38.78 27.52 22.09 25.28 27.26
Baichuan-omni-1.5 27.74 47.22 31.79 28.57 35.78 27.52 29.75 30.32
Ming-lite-omni-1.5 32.85 55.56 37.57 30.61 30.28 29.46 29.98 32.10
Qwen-2.5-omni-7B 40.88 52.78 43.35 38.78 30.28 27.13 29.31 33.23
Gemini-2.0-Flash 46.72 61.11 49.71 63.27 35.78 44.19 44.97 46.29
Gemini-2.5-Flash 67.88 77.78 69.94 85.71 41.28 75.19 67.56 68.23
Gemini-2.5-Pro 81.02 83.33 81.50 91.84 55.05 78.68 74.72 76.61
4 Experiment and Analysis
4.1 Experiment Setting
We evaluate omni-models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-omni(Xu et al., 2025b), Qwen-2.5-omni-3B, Qwen-2.5-omni-7B(Xu et al., 2025a), Baichuan-omni-1.5(Li et al., 2025), MiniCPM-o-2.6(Yao et al., 2024), and Ming-lite-omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). All model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the following sections, we conduct comprehensive experiments on MMAO-Bench and try to answer the following questions:
1. How about the performance of current omni-models and their shortage abilities?
2. What is the relationship of uni-modal ability and omni-modal ability?
3. Could the MMAO-Bench provide an effective "All in One" evaluation for omni-models?
4.2 Model Performance
4.2.1 Overall Analysis
As shown in Table 2, the proprietary Gemini series of models substantially outperform the evaluated open-source counterparts across all benchmarks. Gemini-2.5-Pro consistently leads, establishing a clear state-of-the-art.
When analyzing the cross-modal tasks specifically, we observe distinct patterns. On the Omni-MC (Multiple-choice) benchmark, the performance of smaller open-source models like MiniCPM-o-2.6 (26.77) and Qwen-2.5-omni-3B (27.26) approximates that of random guessing (0.25). In contrast, larger models such as Qwen-2.5-omni-7B (33.23) demonstrate nascent capabilities beyond this baseline. Nevertheless, a significant performance gap remains, with the top open-source model trailing Gemini-2.5-Pro (76.61) by over 43 points.
The Omni-MO (multi-step open-ended) benchmark, which demands more complex reasoning, proves to be even more challenging for all models. Performance universally drops compared to Omni-MC. Most open-source models score poorly, with the best among them, Qwen-2.5-omni-7B, achieving only 29.12. This indicates severe limitations in their ability to perform multi-step cross-modal reasoning. While the Gemini models also see a score reduction, they maintain a commanding lead, with Gemini-2.5-Pro (68.84) showcasing a far more robust, though not perfect, reasoning capability. This highlights that complex, multi-step, cross-modal understanding remains a critical and unresolved challenge, especially for the open-source community.
In terms of perception capabilities, cross-modal alignment is more challenging than cross-modal recognition. Among open-source models, Qwen-2.5-omni-7B has the strongest alignment capability at 40.88, while Ming-lite-omni-1.5 has the strongest recognition capability at 55.56. The Gemini series significantly outperforms other models, and Gemini-2.5-Pro shows no significant weakness in alignment compared to its recognition capability, both exceeding 80.
In terms of reasoning capabilities 3, spatial reasoning is currently the most challenging task, with the best-performing Gemini-2.5-Pro achieving only 55.05. Among open-source models, Baichuan-omni-1.5 performs the best with a score of 35.78. Ming-lite-omni-1.5 demonstrates the best performance in complex reasoning and general reasoning among open-source models.
Overall, perception ability is relatively simpler compared to reasoning ability, and the model gap in reasoning ability is more significant. Taking Qwen-2.5-omni-7B and Gemini-2.5-Pro as examples, they differ by 38.2 points in perception and 45.41 points in reasoning. This demonstrates that reasoning ability is an extension of perception ability.
4.2.2 Top-tier Analysis
Why Gemini-2.5-Pro performances significantly better than other models? We attempt to provide an analysis and some hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, from the perspective of the reasoning process, Gemini possesses audio captioning capabilities, and it can also naturally incorporate audio content as a basis for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, but still has a performance gap on reasoning ability. Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence. The comparison of scores for specific ability items can be seen in Figure6. Dived into ability analysis, we find a interesting phenomena that human is more capable on reasoning than perception (81.3% > 74.3%), and perception ability is even lower than model’s. Through interviewing some of the human experts, we find that human may lost some information in video or audio in one hand, and in the other hand human has limited world knowledge than large language models.
Figure 6: The competition between Human experts and Gemini-2.5-Pro. Gemini-2.5-Pro shows comparable perception capability but lower reasoning capability.
4.3 Uni-Modal v.s. Cross-Modal
To investigate the relationship of uni-modal and omni-modal understanding ability, we conduct regression analysis and ablation experiments. Thanks to the unified ability systems and the high quality of omni-modal samples in MMAO-Bench, we find some interesting observations.
Finding 2. Compositional Law: the performance of omni-modal ability is power-law related to the multiply of uni-modals’ performances. Observing the Table.2, we find that the omni-modal performances are correlation to the multiply of their uni-modals’ performances. In details, the performances of different models are relatively close in audio understanding but show significant differences in vision understanding, and these differences are also pronounced in the omni-modal considering both abilities. Gemini-2.5-Flash and Gemini-2.0-Flash are similar in audio understanding, but for vision modality, Gemini-2.5-Flash outperforms Gemini-2.0-Flash by 9.8, and it is 24.4 stronger in the cross-modal understanding.
Finding 3. The weaker models show "short-board effect", while the stronger models demonstrate emergent ability. We draw the multiply of uni-modal scores as a dot line in the Figure.7. Compared with the cross-modal scores, models around the "random guessing stage" are below the line, while Gemini-2.5-Pro shows a significant gap above the line. Besides, these models demonstrate a power-law tendency that the stronger uni-modal ability can stimulate to much stronger cross-modal ability. This also emphasis that multiple modalities can enhance each other and lead to a higher intelligence.
Next, we conduct ablation studies to dig dive about the enhancement from vision and audio modality respectively.
Figure 7: Relationship between uni-modal ability and cross-modal ability among omni-models.
Table 4: Ablation of visual understanding ability.
Model Perception Reasoning Overall
Audio +Caption +Visual Audio +Caption +Visual Audio +Visual
MiniCPM-o-2.6 26.01 30.06 27.75 24.83 25.28 26.40 25.16 26.77
Qwen-2.5-omni-3B 21.39 29.48 32.37 22.60 24.83 25.28 22.26 27.26
Baichuan-omni-1.5 22.54 30.06 31.79 28.86 31.54 29.75 27.10 30.32
Ming-lite-omni-1.5 26.59 27.17 37.57 25.06 24.38 29.98 25.48 32.10
Qwen-2.5-omni-7B 26.01 33.53 43.35 20.81 25.28 29.31 22.26 33.23
Gemini-2.0-Flash 31.21 53.18 49.71 34.90 51.01 44.97 33.87 46.29
Gemini-2.5-Flash 33.53 68.21 69.94 44.30 70.02 67.56 41.29 68.23
Gemini-2.5-Pro 41.62 78.03 81.50 50.34 73.60 74.72 47.90 76.61
4.3.1 Ablation Visual Understanding
To evaluate the enhancement of visual information on model understanding, we sequentially assessed the performance of Audio (only auditory information), +Caption (auditory + Caption information), and +Visual (auditory information + visual information), where captions were derived from descriptions of videos by Gemini-2.5-Pro.
In the case of audio-only input, most models’ metrics are significantly lower than when both auditory and visual information are provided, and the metrics exhibit a random distribution around 25%. Gemini-2.5-Pro identified problem shortcuts through methods such as elimination, achieving a score of 47.9, which will be optimized in subsequent versions.
After providing the caption information, all model metrics showed significant increases, but the magnitude of the rise varied considerably. In terms of perceptual ability, the Gemini series generally improved by 20 points, while the Qwen-2.5-omni series and Baichuan-omni-1.5 saw an improvement of over 8 points. The MiniCPM-o-2.6 showed a smaller increase of around 4 points, whereas the Ming-lite-omni-1.5 barely improved at all. This indicates that different models have varying abilities to reconstruct image content based on textual descriptions, and similar phenomena are observed in reasoning capabilities as well.
Compared with caption information, the original visual input brings more information. On models with better baseline capabilities, such as Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, caption information is far less than the original visual information. On models with slightly weaker baseline capabilities, the two are approximately the same. Interestingly, the Gemini series has a very high utilization rate of caption information, and even Gemini-2.0-Flash performs better with +caption than with +visual.
4.3.2 Ablation Audio Understanding
We categorize sound types into ambient sounds and dialogue, and provide their transcription versions based on Gemini-2.5-pro, where dialogue is given dialogue text, and ambient sounds are given sound descriptions.
Similar to visual information, in the case of only providing auditory information, most models’ metrics are significantly lower than when both auditory and visual information are provided; after providing speech transcripts, the improvement is again more pronounced in the Gemini series.
MiniCPM-o-2.6 performs better than +Audio on +ASR, confirming that the bottleneck of this model lies in audio understanding.
Table 5: Ablation of audio understanding ability.
Model Perception Reasoning Overall
Video +ASR +Audio Video +ASR +Audio Video +ASR +Audio
MiniCPM-o-2.6 27.75 31.79 27.75 27.07 29.98 26.40 27.26 30.48 26.77
Qwen-2.5-omni-3B 27.17 30.06 32.37 18.79 23.94 25.28 21.13 25.65 27.26
Baichuan-omni-1.5 26.01 27.17 31.79 27.07 23.04 29.75 26.77 24.19 30.32
Ming-lite-omni-1.5 32.37 35.84 37.57 23.04 27.52 29.98 25.65 29.84 32.10
Qwen-2.5-omni-7B 30.06 41.04 43.35 24.16 27.52 29.31 25.81 31.29 33.23
Gemini-2.0-Flash 31.79 55.49 49.71 29.08 47.43 44.97 29.84 49.68 46.29
Gemini-2.5-Flash 38.73 72.83 69.94 37.58 68.68 67.56 37.90 69.84 68.23
Gemini-2.5-Pro 46.82 76.88 81.50 47.20 75.62 74.72 47.10 75.97 76.61
In terms of environmental sound, most open-source models perform poorly. The two more capable models, Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, show better performance in dialogue than in environmental sound. Additionally, we observed that the audio information of MiniCPM-o-2.6 and Ming-lite-omni-1.5 has side effects in environmental sound scenarios. When audio descriptions replace the original audio, both models show significant improvement, indicating that Ming-lite-omni-1.5 and Qwen-2.5-omni-7B have relatively weaker understanding capabilities of audio elements.
In conversational scenarios, the Gemini series achieves ASR results that are nearly on par with, or even surpass, the original audio quality. This distinctive feature is absent in most other open-source models, with the exception of MiniCPM-o-2.6, where the +ASR results exceed +Audio 4 points.
Table 6: Ablation of audio understanding ability.
Model Environment Speech Overall
Visual +Caption +Audio Visual +ASR +Audio Visual +ASR/Caption +Audio
MiniCPM-o-2.6 33.33 31.11 28.89 26.70 30.54 26.53 27.26 30.48 26.77
Qwen-2.5-omni-3B 15.56 31.11 28.89 21.64 25.31 27.23 21.13 25.65 27.26
Baichuan-omni-1.5 26.67 22.22 33.33 26.70 24.43 30.02 26.77 24.19 30.32
Ming-lite-omni-1.5 35.56 37.78 28.89 24.96 29.32 32.29 25.65 29.84 32.10
Qwen-2.5-omni-7B 28.89 40.00 26.67 25.65 30.72 33.86 25.81 31.29 33.23
Gemini-2.0-Flash 37.78 42.22 40.00 29.32 50.44 46.77 29.84 49.68 46.29
Gemini-2.5-Flash 53.33 62.22 51.11 36.47 70.51 69.63 37.90 69.84 68.23
Gemini-2.5-Pro 57.78 71.11 75.56 46.07 76.44 76.79 47.10 75.97 76.61
4.4 Benchmark Analysis
In this section, we verify the effectiveness of MMAO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-sourced benchmarks.
4.4.1 Multi-Step Open-Ended Question Analysis
In this work, we introduce a new type of evaluation method, multi-step open-ended question, which effectively assess the complex reasoning ability, especially appears in cross-modality understanding. As shown in Table.7, the experimental results demonstrate the stratified performance among different models. Gemini-2.5-Pro leads with a comprehensive score of 68.84, showcasing its outstanding capabilities as a top-tier closed-source model. Following closely are Gemini-2.5-Flash and Gemini-2.0-Flash, which also form a clear leading tier. Among open-source models, Qwen-2.5-omni-7B performs best, significantly outperforming other smaller-scale open-source models. This indicates that model scale and advanced architecture remain key factors in addressing complex multi-question tasks.
As the depth of questions increases, model performance generally declines. This is a very significant trend, also validating the effectiveness of our dataset design. For almost all models, their performance on small question 1 is better than on small question 2, while the score for small question 3+ is the lowest. For example, Gemini-2.5-Pro’s score on the "perception" task dropped from 26.60 on small question 1 to 21.79 on small question 3+; on the "reasoning" task, the score sharply decreased from 38.2 on small question 1 to 21.33 on small question 3+. This performance decay phenomenon reveals the common challenges current large models face in handling long-range dependencies, maintaining conversational context, and conducting multi-step chain-of-thought reasoning. Initial perception and simple reasoning are relatively easy, but as conversations deepen, the accumulation of errors and loss of context lead to significantly lower success rates in answering more complex subsequent questions.
Reasoning ability is the key bottleneck that distinguishes model performance. Comparing the total scores of "perception" and "reasoning" dimensions, we found that all models scored higher on perception tasks than on reasoning tasks. For example, Gemini-2.5-Pro’s perception total score is 78.4, while its reasoning total score is 66.45. Among open-source models, Qwen-2.5-omni-7B’s perception total score (40.8) is significantly higher than its reasoning total score (26.20). This indicates that although models have already acquired basic perception capabilities (such as object recognition and sound recognition), there is still significant room for improvement in their ability to transform this perceptual information into complex logical, spatial, or causal reasoning. Our dataset successfully amplifies this capability gap through its progressive question design.
Table 7: Performance on Multi-Step Open-Ended Questions.
Model Perception Reasoning Overall
Q1 Q2 Q3+ All Q1 Q2 Q3+ All Q1 Q2 Q3+ All
Baichuan-omni-3B 8.95 7.75 5.38 24.2 11.8 9.6 9.33 19.55 9.52 8.12 5.87 20.48
MiniCPM-6-2.6B 10.65 9.35 8.11 34.2 22.4 11.8 6.67 18.85 11.16 8.12 6.33 21.92
Qwen-2.5-omni-3B 10.95 8.59 8.76 34.2 19.4 13.2 5.33 23.60 12.64 9.52 8.33 25.72
Ming-lite-omni 8.35 7.20 6.29 33.2 20.8 12.4 4.67 24.30 12.68 9.96 7.69 26.08
Qwen-2.5-omni-7B 12.85 9.75 7.90 40.8 18.4 17.6 16.00 26.20 13.96 11.33 8.92 29.12
Gemini-2.0-Flash 17.55 13.60 11.32 56.2 27.8 22.0 21.33 37.15 19.6 15.28 12.56 40.96
Gemini-2.5-Flash 23.40 25.05 20.66 72.4 35.2 31.4 26.00 59.40 25.76 26.32 21.32 62.00
Gemini-2.5-Pro 26.60 29.10 21.79 78.4 38.2 33.8 21.33 66.45 28.92 30.04 21.74 68.84
4.4.2 Dataset Compression
We design a cluster-guided stratified sampling to compress the scale of benchmark. To evaluate the consistency of model ranking and the best size of compression data size, we conduct several experiments to analysis.
Baseline dataset consists of  8000 samples including 15 open-source benchmarks (e.g. MathVista and MMAU, details see appendix) and 12 models evaluation results on them, which splits into 8/4 on models as training/test set. Kmeans is used to cluster with K=48. To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
To verify the compression performance, we use metrics Margin of Error(MoE), Root Mean Square Error(RMSE), CI Coverage, Spearman’s linear correlation coefficient(SRCC) and Pearson’s linear correlation coefficient(PLCC). The experimental result is shown in Figure.8, when compression rate is 10%, the coverage is nearly 100% and PLCC is above 0.98, SRCC is above 0.98, RMSE is 0.06, which means the model ranking is consistent.
Figure 8: Data compression Performance.
4.4.3 Benchmark Comparison
Besides dataset quality and evaluation efficiency, a good benchmark should provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth. Compared with existing omni-modal benchmarks (including OmniBench(Li et al., 2024b), WorldSense(Hong et al., 2025) and AVOdyssey(Gong et al., 2024)), MMAO-Bench has relatively consistent rankings but a higher standard deviation of 19.5 in evaluation scores, while the best other OmniBench only has 12.75, showing the MMAO-Bench has greater discriminative ability. A simple way to improve discrimination is increasing problem difficulty, which can’t guide model’s growth. As a verification, we can see that Gemini-2.5-pro scored 76.61 on MMAO-Bench, but only 65.49 on OmniBench, indicating that MMAO-Bench has appropriate difficulty levels.
Figure 9: Comparison between omni benchmarks.
5 Conclusion
In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively. The main foundings are summarized as below.
1. With unified data framework in MMAO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
2. Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni-models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
3. Besides better dataset quality and evaluation efficiency, MMAO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
In the future work, we will extend the dataset’s scale by human-in-the-loop automatic pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore.